if i have 420 what percentage would i have to take away to get 415?
1
Expert's answer
2013-06-25T12:05:01-0400
At first, we need to know what persentage of 420 is 415:
415/420*100%=approximately 0,988*100%=98,8%
So 415 is equal to 98,8% of 420. Thus, we need to take away 100-98,8=1,2% of 420 to get 415.